This year John's mother is exactly three times as old as he is. Last year, their ages added up to 50. How old is John?
5x3 = 45, John's mother must be 45 and John is 5 5+45 = 50 is that correct?

Answer by mananth(16949)

Let John’s age be x
Mothers age = 3x
Last year john’s age = x-1
Mother’s age last year = 3x-1
The sum of their ages last year was 50
x-1 + 3x-1 = 50
4x-2=50
4x=52
X=13 = John’s age


M Ananth___________ mananth@hotmail.com